OpenStudy (anonymous):

Marks,jasper and Eric are brothers. Mark is 2 years less than twice the age of Eric. Jasper is four years younger than eric. the sum of their ages is 22. write an equation in one variable that can be used to find eric's age. Determine the ages of mark and jasper. @perl

OpenStudy (perl):

M = 2*E - 2 J= E - 4 M + J + E = 22

OpenStudy (anonymous):

What do i do first @perl ?

OpenStudy (perl):

from the second equation you can substitute E in terms of J

OpenStudy (perl):

ola, you can also plug it into wolfram like this wolframalpha.com "solve M = 2*E - 2, J= E - 4, M + J + E = 22"

OpenStudy (anonymous):

yeah @perl

OpenStudy (perl):

@Miracrown can you look this over

OpenStudy (anonymous):

Mark,Jasper and Eric are Brothers. Mark is two years less than twice the age of eric. Jasper is 4 years younger than eric. the sum of their ages is 22.

OpenStudy (perl):

ill try to find someone to check this over. its possible i made a mistake somewhere

OpenStudy (perl):

@gorv can you check this . did i read this wrong

OpenStudy (gorv):

pug in the values in M+J+E=22

OpenStudy (gorv):

2E-2+E-4+E=22 4E-6=22

OpenStudy (perl):

M = 12,E = 7, J = 3

OpenStudy (perl):

ok for some reason wolfram said there is no solution

OpenStudy (gorv):

yeah bro :)

OpenStudy (gorv):

u can match the value

OpenStudy (gorv):

M=2*7-2=12

OpenStudy (anonymous):

Thank you guys i was really stuck !! @perl @gorv

OpenStudy (gorv):

no lol just remove the solve from there

OpenStudy (gorv):

@perl u r smart mathematician ..nobody can let u down :)

OpenStudy (perl):

yeah, but do you know the command to solve a system of equations for wolfram?

OpenStudy (perl):

theres no solution there, the lines do not intersect

OpenStudy (gorv):

no lol i used by hand alwys ..it was just trial and error....i tried removing solve

OpenStudy (perl):

oh, for some reason wolfram wants to force these into lines. they are intersections of 3 planes

OpenStudy (perl):

not three lines

OpenStudy (gorv):

i agree ..i also think so ..well prob is done :)

OpenStudy (perl):

thankyou gorv for your comments and insight

OpenStudy (perl):

that was clever to remove 'solve', now i see that wolfram thought it was three lines

OpenStudy (perl):

i have an idea, lets use x,y,z

OpenStudy (gorv):

lol no need to thank....we are helpin others here ...doin good ....u would have done the same if i have asked lol

OpenStudy (gorv):

that will be better

OpenStudy (perl):

it worked now

OpenStudy (perl):

computers are finicky :)

OpenStudy (gorv):

hahhaha :D